### v3

# MolFORM: Preference-Aligned Multimodal Flow Matching for Structure-Based Drug Design

Abstract.
Structure-based drug design (SBDD) aims to efficiently discover high-affinity ligands within vast chemical spaces. However, current generative models struggle with objective misalignment and rigid sampling budgets. We present MolFORM, a fast multi-modal flow matching framework for discrete atom types and continuous coordinates. Crucially, to bridge the gap between generative capability and biochemical objectives, we introduce two distinct post-training strategies: (1) Direct Preference Optimization (DPO), which performs offline alignment using ranked preference pairs; and (2) an online reinforcement learning paradigm that optimizes the generative flow directly on the forward process. Both strategies effectively navigate the chemical space toward high-affinity regions. MolFORM achieves state-of-the-art results on CrossDocked2020 benchmark (Vina Score -7.60, Diversity 0.75), demonstrating that incorporating preference alignment mechanisms—whether via offline optimization or online reinforcement—is crucial for steering generative models toward high-affinity binding regions. The source code for MolFORM is publicly available at https://github.com/daiheng-zhang/SBDD-MolFORM.
Structure based drug design, Flow matching, Preference alignment
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Structure-based drug design (SBDD) (Anderson, 2003) accelerates drug discovery by utilizing the three-dimensional structures of biological targets, enabling the efficient and rational design of molecules within a defined chemical space. Generative models have recently emerged as a powerful approach for streamlining the SBDD process by directly proposing candidate molecules, thus bypassing the need for exhaustive exploration of large chemical libraries. Advances in this area can be broadly categorized into two directions: autoregressive models  (Luo et al., 2021), which formulate molecule generation as a sequential prediction task, and diffusion models  (Guan et al., 2023, 2024), which draw inspiration from the iterative refinement process commonly used in image generation. Despite the variety of non-autoregressive generative models, diffusion-based approaches have become the dominant paradigm. In SBDD, many extensions have been developed on top of diffusion models to better handle protein-ligand interactions, with a particular focus on improving binding affinity through task-specific objectives and interaction-aware designs  (Huang et al., 2023; Guan et al., 2024). In parallel, there has been growing interest in exploring alternative non-autoregressive frameworks such as Bayesian Flow Networks  (Qu et al., 2024), which have also demonstrated promising results, achieving state-of-the-art performance  (Lin et al., 2024) on several benchmark SBDD tasks.
In recent years, flow matching (Liu et al., 2022b; Lipman et al., 2022) has emerged as a widely studied generative modeling framework. Although flow matching is theoretically equivalent to diffusion models under certain conditions (Gao et al., 2024), empirical performance can vary significantly depending on the choice of scheduling strategy. In image generation tasks, flow matching has been successfully scaled to large datasets and has demonstrated strong performance (Esser et al., 2024; Liu et al., 2023). In the domain of AI for science, especially for molecular and protein generation tasks, researchers have also begun to explore the applicability of flow matching (Jing et al., 2024; Geffner et al., 2025; Campbell et al., 2024). Conceptually, flow matching offers a transport mapping interpretation from the perspective of ordinary differential equations (ODEs), providing a flexible modeling framework that enables task-specific adaptations. For structure-based drug design (SBDD), the generative task involves predicting both atom types and their 3D positions, which can be viewed as a combination of discrete and continuous modalities. Motivated by recent advances, we propose MolFORM, a novel framework for Molecular multi-modal Flow-Optimized Representation Matching. To further enhance sampling efficiency, we also designed an auxiliary confidence head capable of predicting confidence scores for generated structures, serving as a basis for ranking high-quality candidates.
Furthermore, to bridge the gap between generative capability and biochemical objectives, we introduce the preference-guided fine-tuning stage comprising two distinct strategies: offline Direct Preference Optimization (DPO) (Rafailov et al., 2023; Wallace et al., 2023) and online Reinforcement Learning (RL) (Zheng et al., 2025). We demonstrate that the substantial performance gains stem from a multi-flow co-modeling strategy that jointly aligns preferences over both discrete atom identities and continuous 3D positions. Beyond offline alignment, we further incorporate an online RL paradigm that optimizes the generative flow directly on the forward process. By dynamically contrasting positive and negative generations sampled during training, this approach efficiently navigates the chemical space toward high-affinity regions. Our experiments show that these rigorous flow alignment techniques leverage the Vina score (Trott and Olson, 2010) as a chemically informed reward signal to significantly enhance molecule quality, while enabling the flexible design of reward functions tailored to specific requirements for the target molecule.
2. Related work
2.1. Structure-Based Drug Design.
With the increasing availability of structural data, generative models have attracted significant attention for structure-based molecule generation. Early methods (Skalic et al., 2019) utilized sequence generative models to produce SMILES representations from protein contexts. Driven by advancements in 3D geometric modeling, subsequent studies directly generate molecules in 3D space. For instance, Ragoza et al. (2022) employ voxelized atomic density grids within a Variational Autoencoder framework. Other methods use autoregressive models to sequentially place atoms or chemical groups (Luo et al., 2021; Peng et al., 2022), while FLAG (Zhang et al., 2023) and DrugGPS (Zhang and Liu, 2023) leverage chemical priors to construct realistic ligand fragments incrementally. More recently, diffusion models have demonstrated notable success by progressively denoising atom types and coordinates, maintaining SE(3)-equivariant symmetries (Guan et al., 2023; Lin et al., 2022; Schneuing et al., 2023; Huang et al., 2023; Guan et al., 2024; Zhang et al., 2024). Despite these advances, existing models often struggle with generating molecules simultaneously optimized for multiple desirable properties, such as binding affinity, synthesizability, and low toxicity in drug discovery (D Segall, 2012).
2.2. Flow Matching.
Flow matching (Lipman et al., 2022; Liu et al., 2022b) is a continuous-time generative modeling framework that generalizes diffusion models by learning a time-dependent vector field to transport a simple prior $q(\bm{x})$ toward the data distribution $p_{\text{data}}(\bm{x})$. It defines a conditional probability path $p_{t}(\bm{x}\mid\bm{x}_{1})$ that interpolates between $q(\bm{x})$ and the target $\delta(\bm{x}-\bm{x}_{1})$, and learns the corresponding marginal vector field $v(\bm{x},t)=\mathbb{E}_{\bm{x}_{1}\sim p_{t}(\bm{x}_{1}\mid\bm{x})}[u_{t}(\bm{x}\mid\bm{x}_{1})]$ using a neural network $v_{\theta}(\bm{x},t)$. Diffusion models can be viewed as a special case of flow matching under Gaussian interpolation (Gao et al., 2024). As a simulation-free training paradigm for continuous normalizing flows, flow matching offers flexibility in choosing probability paths and time schedules, while recent method such as Rectified Flow encourage straighter transport trajectories and enable efficient sampling with fewer ODE steps. Flow matching has demonstrated strong performance in large-scale image and video generation (Liu et al., 2023; Esser et al., 2024), and is increasingly adopted in scientific domains such as protein generation (Campbell et al., 2024) and protein conformation modeling (Jing et al., 2024). Moreover, discrete flow models extend flow matching to discrete state spaces via continuous-time Markov chains, recovering discrete diffusion as a special case and enabling unified modeling of multimodal settings that couple discrete atom types with continuous 3D structures.
2.3. Preference Alignment of Diffusion Models.
While maximizing data likelihood is standard in generative modeling, it often fails to align with downstream user preferences. Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2020; Ouyang et al., 2022) has been widely adopted to align large language models with human intent. Recent efforts extend these ideas to diffusion models by treating generation as a multi-step decision process (Uehara et al., 2024; Wallace et al., 2023). To implement this, early works typically formulate sampling as a Markov Decision Process (MDP), discretizing the reverse process to apply Policy Gradient algorithms (Black et al., 2023; Fan et al., 2023). Notably, methods like FlowGRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025) successfully adapt Group Relative Policy Optimization (GRPO) to diffusion by leveraging SDE-based stochasticity for exploration. However, these reverse-process approaches often suffer from solver restrictions and forward-reverse inconsistency. To address these limitations, DiffusionNFT (Zheng et al., 2025) introduces a novel online RL paradigm that optimizes directly on the forward process via flow matching. By contrasting positive and negative generations to define an implicit improvement direction, DiffusionNFT eliminates the need for likelihood estimation and achieves significantly higher training efficiency compared to GRPO-based methods.
Alternatively, Direct Preference Optimization (DPO) (Rafailov et al., 2023) offers a simpler paradigm by bypassing reinforcement learning and directly optimizing models against pairwise preference data. DPO has shown competitive results in both language and image domains (Wallace et al., 2023; Zhou et al., 2024b). In the context of structure-based drug design (SBDD), recent studies (Gu et al., 2024; Cheng et al., 2024) have also begun to incorporate preference alignment to improve biological plausibility and design success.
3. Methods
3.1. Problem definitions.
We aim to generate ligand molecules that are capable of binding to specific protein binding sites, by modeling $p(M|P)$. We represent the protein pocket as a collection of $N_{P}$ atoms, $P=\{(x^{i}_{P},v^{i}_{P})\}_{i=1}^{N_{P}}$. Similarly, the ligand molecule can be represented as a collection of $N_{M}$ atoms, $M=\{(x^{i}_{M},v^{i}_{M})\}_{i=1}^{N_{M}}$, where $x^{i}_{M}\in\mathbb{R}^{3}$ represents atom position and $v^{i}_{M}\in[k]$ the $k$ possible atom types. The number of atoms $N_{M}$ can be sampled from an empirical distribution (Hoogeboom et al., 2022; Guan et al., 2023). For brevity, the ligand molecule is denoted as $M=\{\mathbf{X},\mathbf{V}\}$ where $\mathbf{X}\in\mathbb{R}^{N_{M}\times 3}$ and $\mathbf{V}\in[k]^{N_{M}\times{K}}$.
Figure 1. Overview of MolFORM. This workflow can be summarized as two steps: 1) Employs multi-flow generation to construct the base model. 2) Applies DPO to fine-tune the dual modalities, using the Vina score as the reward.
3.2. Multi-modal Flow Matching
The overall framework of MolFORM is illustrated in Figure 1. We model a ligand as a multimodal object consisting of continuous atomic coordinates $x\in\mathbb{R}^{N_{M}\times 3}$ and discrete atom types $v\in[K]^{N_{M}}$, conditioned on a protein pocket $p$. MolFORM jointly learns (i) a continuous flow for coordinates and (ii) a discrete flow for atom types, using a shared SE(3)-equivariant backbone and synchronized time $t\in[0,1]$.
Continuous flow matching for atomic coordinates.
Conditional Flow Matching (CFM) learns a time-dependent flow $\psi_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that transports samples from a source distribution $p_{0}$ to a target distribution $p_{1}$, governed by the ODE $\frac{d}{dt}x_{t}=u_{t}(x_{t})$ with $x_{t}=\psi_{t}(x_{0})$. Since the exact marginal vector field $u_{t}$ is generally intractable, CFM defines a conditional vector field $u_{t}(x_{t}\mid x_{0},x_{1})$ and trains a neural vector field (velocity) $\mathbf{v}_{\theta}$ by regression:
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{t,\,x_{0},\,x_{1}}\left\|\mathbf{v}_{\theta}(x_{t},v_{t},p,t)-u_{t}(x_{t}\mid x_{0},x_{1})\right\|_{2}^{2},$
where we omit conditioning variables in $u_{t}(\cdot)$ for brevity.
In this work, we use the rectified flow on Euclidean space:
$x_{t}=(1-t)x_{0}+tx_{1},\qquad u_{t}(x_{t}\mid x_{0},x_{1})=x_{1}-x_{0},$
with $x_{0}\sim p_{0}$ a Gaussian prior and $x_{1}\sim p_{\mathrm{data}}$ a data sample.
Discrete flow matching for atom types.
For discrete atom types, we adopt the Discrete Flow Matching based on continuous-time Markov chains (CTMCs). We define a family of conditional flows $\pi_{t}(v_{t}\mid v_{1})$ and use uniform corruption:
$\pi_{t}(v_{t}\mid v_{1})=(1-t)\cdot\mathrm{Uniform}([K])+t\cdot\delta_{v_{1}}(v_{t}),$
where $\delta_{v_{1}}(v_{t})$ is the Kronecker delta and $\mathrm{Uniform}([K])=1/K$. The corresponding marginal at time $t$ is
$p_{t}(v_{t})=\mathbb{E}_{v_{1}\sim p_{\mathrm{data}}}\big[\pi_{t}(v_{t}\mid v_{1})\big].$
We train a denoising posterior $p_{\theta,1|t}(v_{1}\mid v_{t},p,t)$ to predict clean atom types. A standard cross-entropy objective is
$\mathcal{L}_{\mathrm{CE}}=\mathbb{E}_{t,\,v_{1},\,v_{t}\sim\pi_{t}(\cdot\mid v_{1})}\left[-\sum_{i=1}^{N_{M}}\log p_{\theta,1|t}\!\left(v^{i}_{1}\mid v_{t},p,t\right)\right],$
where $v_{1}=\{v_{1}^{i}\}_{i=1}^{N_{M}}$ and $v_{t}=\{v_{t}^{i}\}_{i=1}^{N_{M}}$.
Reparameterized training with $x$-prediction.
Although $\mathbf{v}_{\theta}$ is the velocity field used for sampling, we adopt a reparameterized training objective that computes regression in $x$-space for improved numerical stability. Given the current noisy state $(x_{t},v_{t})$, we form a one-step endpoint reconstruction from time $t$ to $1$:
$\hat{x}_{1}\;=\;x_{t}+(1-t)\,\mathbf{v}_{\theta}(x_{t},v_{t},p,t).$
We then define the reconstructed-flow loss
$\mathcal{L}_{\mathrm{reparam}}(\theta)=\mathbb{E}_{t,\,x_{0},\,x_{1}}\left\|u_{t}(x_{t}\mid\hat{x}_{1},x_{0})-u_{t}(x_{t}\mid x_{1},x_{0})\right\|_{2}^{2}.$
On Euclidean manifolds with the straight-line path $u_{t}(x_{t}\mid x_{0},x_{1})=x_{1}-x_{0}$, this reduces to a simple $x$-space MSE:
$\mathcal{L}_{\mathrm{pos}}=\mathbb{E}_{t,\,x_{0},\,x_{1}}\left\|\hat{x}_{1}-x_{1}\right\|_{2}^{2}.$
Similarly, for discrete atom types we directly predict the clean posterior $p_{\theta,1|t}(\cdot\mid v_{t},p,t)$ and use the corresponding cross-entropy (this fixes the $v_{0}$$v_{t}$ notation issue):
$\mathcal{L}_{\mathrm{type}}=\mathbb{E}_{t,\,v_{1},\,v_{t}\sim\pi_{t}(\cdot\mid v_{1})}\left[\mathrm{CE}\!\left(p_{\theta,1|t}(\cdot\mid v_{t},p,t),\,v_{1}\right)\right].$
Chamfer loss.
To promote accurate geometric alignment between predicted and ground-truth molecular structures, we incorporate a Chamfer loss defined over atomic point clouds. Given two point sets $\hat{x}_{1}=\{\hat{x}_{i}\}_{i=1}^{N}$ and $x_{1}=\{x_{j}\}_{j=1}^{M}$ representing predicted and reference atomic positions respectively, the Chamfer distance is
(1) $\mathcal{L}_{\mathrm{Chamfer}}=\frac{1}{N}\sum_{\hat{x}\in\hat{x}_{1}}\min_{x\in x_{1}}\|\hat{x}-x\|_{2}+\frac{1}{M}\sum_{x\in x_{1}}\min_{\hat{x}\in\hat{x}_{1}}\|x-\hat{x}\|_{2}.$
Overall objective.
The final pretraining objective for the multimodal flow is
$\mathcal{L}=\mathcal{L}_{\mathrm{pos}}+\mathcal{L}_{\mathrm{type}}+\lambda\cdot\mathcal{L}_{\mathrm{Chamfer}},$
where $\lambda$ is a weighting hyperparameter.
3.3. Sampling
We consider two generative sampling procedures, the continuous atomic coordinates and discrete atom types. Starting from a joint noise prior $(x_{0},v_{0})$ with $x_{0}\sim p_{0}$ (Gaussian) and $v_{0}\sim\mathrm{Uniform}([K])$, evolving both modalities forward from $t=0$ to $t=1$.
Continuous sampling.
For atomic coordinates, we simulate trajectories via the learned velocity field $\mathbf{v}_{\theta}(x_{t},v_{t},p,t)$ using Euler integration in $N$ steps:
(2) $x_{t+\frac{1}{N}}=x_{t}+\frac{1}{N}\cdot\mathbf{v}_{\theta}(x_{t},v_{t},p,t),\qquad t\in\left\{0,\frac{1}{N},\ldots,\frac{N-1}{N}\right\}.$
Integrating this ODE from $x_{0}$ yields the final conformation $x_{1}$.
Discrete sampling.
For atom types, we simulate a CTMC using an Euler discretization. Given the current type $v_{t}$,
$v_{t+\Delta t}\sim\mathrm{Cat}\!\left(\delta_{v_{t}}+R_{\theta,t}(v_{t},\cdot\mid p)\,\Delta t\right),$
where $R_{\theta,t}$ is the model-induced (unconditional) rate matrix and $\delta_{v_{t}}$ denotes the one-hot vector at $v_{t}$. The unconditional rate can be computed as a posterior expectation over the forward conditional rate matrix $R^{q}_{t}(\cdot,\cdot\mid v_{1})$:
$R_{\theta,t}(v_{t},j\mid p)=\mathbb{E}_{v_{1}\sim p_{\theta,1|t}(\cdot\mid v_{t},p,t)}\left[R^{q}_{t}(v_{t},j\mid v_{1})\right].$
The exact Bayesian posterior satisfies
$q(v_{1}\mid v_{t})=\frac{\pi_{t}(v_{t}\mid v_{1})\,p_{\mathrm{data}}(v_{1})}{p_{t}(v_{t})},$
and in practice we approximate $q(v_{1}\mid v_{t})$ using the learned denoising posterior $p_{\theta,1|t}(v_{1}\mid v_{t},p,t)$. (Under the uniform corruption in Eq. (1), this also yields the convenient closed form $R_{\theta,t}(v_{t},j\mid p)=\frac{1}{1-t}\,p_{\theta,1|t}(v_{1}=j\mid v_{t},p,t)$ for $j\neq v_{t}$.
3.4. Direct Preference Optimization (DPO) on Multi-Flow
Multi-flow model provides strong generative capability, downstream objectives (e.g., docking affinity) are not guaranteed to be aligned with the model likelihood. We therefore perform an offline preference-alignment stage using Direct Preference Optimization (DPO). For each protein pocket condition $p$, we collect a winner molecule and a loser molecule and form a preference pair. A molecule at the terminal time is denoted by $m_{1}:=(x_{1},v_{1})$, where $x_{1}$ are continuous atomic coordinates and $v_{1}$ are discrete atom types. The preference dataset is $\mathcal{D}_{\mathrm{pref}}=\{(p,m_{1}^{w},m_{1}^{l})\}$.
Standard DPO objective.
Let $p_{\theta}(m_{1}\mid p)$ be the learnable conditional generative model and $p_{\mathrm{ref}}(m_{1}\mid p)$ be a fixed reference (the pretrained base model). DPO optimizes $\theta$ via a binary classification objective:
$\mathcal{L}_{\mathrm{DPO}}(\theta)$ $=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}}}\Big[\log\sigma\!\Big(\beta\big(\log\tfrac{p_{\theta}(m_{1}^{w}\mid p)}{p_{\mathrm{ref}}(m_{1}^{w}\mid p)}$
$\qquad-\log\tfrac{p_{\theta}(m_{1}^{l}\mid p)}{p_{\mathrm{ref}}(m_{1}^{l}\mid p)}\big)\Big)\Big].$
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the logistic sigmoid and $\beta>0$ controls the strength of preference regularization.
Flow-based surrogate for multimodal data.
For flow matching models, directly evaluating $\log p_{\theta}(m_{1}\mid p)$ is intractable. Following the common practice in diffusion/flow preference alignment, we instead define a timestep-wise surrogate by sampling $t\sim\mathcal{U}[0,1]$ and corrupting each modality with the same forward processes used in pretraining:
$x_{t}\sim q_{t\mid 1}(x_{t}\mid x_{1}),\qquad v_{t}\sim\pi_{t}(v_{t}\mid v_{1}).$
The model predicts the clean sample at time $t$: $\hat{x}_{1,\theta}=\hat{x}_{1,\theta}(x_{t},v_{t},p,t)$ (and analogously $\hat{x}_{1,\mathrm{ref}}$ from the reference model), and predicts a categorical reverse posterior $p_{\theta,1\mid t}(v_{1}\mid v_{t},p)$ (and $p_{\mathrm{ref},1\mid t}$).
To align both continuous coordinates and discrete atom types, we apply DPO separately on each modality and optimize their weighted sum (consistent with the base multi-flow training loss structure). Below we write three modality-wise DPO losses.
DPO on continuous coordinates.
$\mathcal{L}_{\mathrm{DPO}}^{x}(\theta)=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}},\;t\sim\mathcal{U}[0,1]}\Big[\log\sigma\Big(\beta\big(\Delta_{x}^{w}-\Delta_{x}^{l}\big)\Big)\Big],$
where
$\Delta_{x}^{w}$ $=-\big\|x_{1}^{w}-\hat{x}_{1,\theta}^{w}\big\|_{2}^{2}+\big\|x_{1}^{w}-\hat{x}_{1,\mathrm{ref}}^{w}\big\|_{2}^{2},$
$\Delta_{x}^{l}$ $=-\big\|x_{1}^{l}-\hat{x}_{1,\theta}^{l}\big\|_{2}^{2}+\big\|x_{1}^{l}-\hat{x}_{1,\mathrm{ref}}^{l}\big\|_{2}^{2},$
and the noisy states are sampled as $x_{t}^{w}\!\sim q_{t\mid 1}(\cdot\mid x_{1}^{w})$, $x_{t}^{l}\!\sim q_{t\mid 1}(\cdot\mid x_{1}^{l})$, $v_{t}^{w}\!\sim\pi_{t}(\cdot\mid v_{1}^{w})$, $v_{t}^{l}\!\sim\pi_{t}(\cdot\mid v_{1}^{l})$, with $\hat{x}_{1,\theta}^{w}=\hat{x}_{1,\theta}(x_{t}^{w},v_{t}^{w},p,t)$ and $\hat{x}_{1,\theta}^{l}=\hat{x}_{1,\theta}(x_{t}^{l},v_{t}^{l},p,t)$ (and similarly for the reference model).
DPO on point-cloud geometry (Chamfer).
To additionally enforce geometric alignment at the point-cloud level, we apply the same preference objective to the Chamfer distance $L_{\mathrm{Chamfer}}(\cdot,\cdot)$ defined in Eq. 1:
$\mathcal{L}_{\mathrm{DPO}}^{\mathrm{pc}}(\theta)=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}},\;t\sim\mathcal{U}[0,1]}\Big[\log\sigma\Big(\beta\big(\Delta_{\mathrm{pc}}^{w}-\Delta_{\mathrm{pc}}^{l}\big)\Big)\Big],$
where
$\Delta_{\mathrm{pc}}^{w}$ $=-L_{\mathrm{Chamfer}}\!\big(x_{1}^{w},\hat{x}_{1,\theta}^{w}\big)+L_{\mathrm{Chamfer}}\!\big(x_{1}^{w},\hat{x}_{1,\mathrm{ref}}^{w}\big),$
$\Delta_{\mathrm{pc}}^{l}$ $=-L_{\mathrm{Chamfer}}\!\big(x_{1}^{l},\hat{x}_{1,\theta}^{l}\big)+L_{\mathrm{Chamfer}}\!\big(x_{1}^{l},\hat{x}_{1,\mathrm{ref}}^{l}\big).$
DPO on discrete atom types via CTMC rates.
For discrete flow matching, we follow the CTMC rate-matrix parameterization and define a relative rate-based surrogate between the current model and the reference model:
$\mathcal{L}_{\mathrm{DPO}}^{v}(\theta)=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}},\;t\sim\mathcal{U}[0,1]}\Big[\log\sigma\Big(\beta\big(\mathcal{D}^{\theta}_{\mathrm{ref}}(v_{t}^{w}\mid v_{1}^{w},p,t)\Big$.
$\Big.-\mathcal{D}^{\theta}_{\mathrm{ref}}(v_{t}^{l}\mid v_{1}^{l},p,t)\big)\Big)\Big].$
The relative rate score $\mathcal{D}^{\theta}_{\mathrm{ref}}$ compares the model-induced unconditional rate matrix $R_{t}^{\theta}(\cdot,\cdot\mid p)$ with the reference rate matrix $R_{t}^{\mathrm{ref}}(\cdot,\cdot\mid p)$ under the forward conditional rate $R_{t}^{q}(\cdot,\cdot\mid v_{1})$:
$\mathcal{D}^{\theta}_{\mathrm{ref}}(v_{t}\mid v_{1},p,t)$ $=\sum_{j\neq v_{t}}\left[R_{t}^{q}(v_{t},j\mid v_{1})\,\log\frac{R_{t}^{\theta}(v_{t},j\mid p)}{R_{t}^{\mathrm{ref}}(v_{t},j\mid p)}\right$.
(3) $\left.\quad+R_{t}^{\mathrm{ref}}(v_{t},j\mid p)-R_{t}^{\theta}(v_{t},j\mid p)\right].$
The model-induced unconditional rate matrix is the posterior expectation of the conditional rate:
$R_{t}^{\theta}(v_{t},j\mid p)=\mathbb{E}_{v_{1}\sim p_{\theta,1\mid t}(\cdot\mid v_{t},p)}\left[R_{t}^{q}(v_{t},j\mid v_{1})\right],\qquad j\neq v_{t},$
and similarly for $R_{t}^{\mathrm{ref}}$ using $p_{\mathrm{ref},1\mid t}$. (As usual in CTMCs, the diagonal is $R_{t}^{\theta}(v_{t},v_{t}\mid p)=-\sum_{j\neq v_{t}}R_{t}^{\theta}(v_{t},j\mid p)$.)
Uniform noising simplification.
Under the uniform corruption initialization $\mathrm{Uniform}([k])$ for the discrete forward process, the unconditional rate admits a closed form:
$R_{t}^{\theta}(v_{t},j\mid p)$ $=\frac{1}{1-t}\,p_{\theta,1\mid t}(v_{1}=j\mid v_{t},p),$
(4) $R_{t}^{\mathrm{ref}}(v_{t},j\mid p)$ $=\frac{1}{1-t}\,p_{\mathrm{ref},1\mid t}(v_{1}=j\mid v_{t},p),\qquad j\neq v_{t}.$
Substituting Eq. (4) into Eq. (3) yields a discrete DPO surrogate expressed purely in terms of the reverse categorical posteriors:
$\mathcal{L}_{\mathrm{DPO}}^{v}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}}\\ t\sim\mathcal{U}[0,1]\end{subarray}}\Bigg[\log\sigma\Bigg(\frac{\beta}{1-t}\Big(\log\frac{p_{\theta,1\mid t}(v_{1}^{w}\mid v_{t}^{w},p)}{p_{\mathrm{ref},1\mid t}(v_{1}^{w}\mid v_{t}^{w},p)}$
$\qquad-\log\frac{p_{\theta,1\mid t}(v_{1}^{l}\mid v_{t}^{l},p)}{p_{\mathrm{ref},1\mid t}(v_{1}^{l}\mid v_{t}^{l},p)}\Big)\Bigg)\Bigg].$
We jointly fine-tune the multi-flow model:
$\mathcal{L}_{\mathrm{MF\text{-}DPO}}(\theta)=\mathcal{L}_{\mathrm{DPO}}^{x}(\theta)+\lambda\,\mathcal{L}_{\mathrm{DPO}}^{\mathrm{pc}}(\theta)+\mathcal{L}_{\mathrm{DPO}}^{v}(\theta),$
where $\lambda$ is the same geometry weighting coefficient.
Figure 2. Online RL on Multi-Flow. By implicitly parameterizing positive and negative branches relative to the reference policy (instead of learning a separate guidance model), we integrate reinforcement guidance directly into the flow-matching objective and steer the generative flow toward high-reward regions.
3.5. Online RL on Multi-Flow
Problem Setup.
In the online reinforcement learning stage, we maintain an anchor (sampling) policy $\pi_{\theta_{\text{old}}}(\cdot\mid p)$ over complete ligand structures $M=(x_{1},v_{1})$ conditioned on a protein pocket $p$. At each iteration, we roll out $K$ candidate molecules $M_{i}\sim\pi_{\theta_{\text{old}}}(\cdot\mid p)$ and evaluate each sample by a biochemical reward $R_{\text{raw}}(M_{i},p)$. Throughout this section, we assume $R_{\text{raw}}$ is higher-is-better; for docking energies such as Vina (lower-is-better), we use a monotone transform (e.g., $\text{Vina}$) as $R_{\text{raw}}$. The complete procedure is summarized in Algorithm 1.
We introduce a latent binary optimality variable $o\in\{0,1\}$, where $o=1$ denotes a high-quality molecule. We define an optimality probability
$r(M,p):=\mathbb{P}(o=1\mid M,p)\in[0,1],$
obtained from a normalized reward transformation (detailed below). The marginal optimality under the anchor policy is
$\mathbb{P}(o=1\mid p)=\mathbb{E}_{M\sim\pi_{\theta_{\text{old}}}(\cdot\mid p)}\big[r(M,p)\big].$
By the law of total probability, the anchor policy admits a mixture decomposition:
$\pi_{\theta_{\text{old}}}(M\mid p)=\mathbb{P}(o=1\mid p)\,\pi^{+}(M\mid p)+\mathbb{P}(o=0\mid p)\,\pi^{-}(M\mid p),$
where $\pi^{+}(M\mid p):=\pi(M\mid o=1,p)$ is the high-reward component and $\pi^{-}(M\mid p):=\pi(M\mid o=0,p)$ is the low-reward component. Our goal is to improve the policy by steering probability mass toward $\pi^{+}(\cdot\mid p)$ while reducing mass on $\pi^{-}(\cdot\mid p)$, without explicit likelihood-ratio estimation.
Reinforcement Guidance Direction.
Following DiffusionNFT, we interpret $r(M,p)$ as a soft indicator that implicitly partitions rollouts from $\pi_{\theta_{\text{old}}}(\cdot\mid p)$ into positive and negative components. Instead of applying policy gradients on intractable likelihoods, DiffusionNFT defines a guided target vector field for the continuous flow:
$v^{*}(x_{t},v_{t},p,t)=v_{\text{old}}(x_{t},v_{t},p,t)+\frac{1}{\beta}\,\Delta(x_{t},v_{t},p,t),$
where $\Delta$ is a reinforcement improvement direction and $\beta>0$ controls the trade-off between staying close to the anchor policy and moving along the improvement direction (the effective guidance strength scales with $1/\beta$). Since $\pi^{+}$ and $\pi^{-}$ are implicit and intractable, we do not explicitly estimate them. Instead, we realize this guidance through an implicit positive/negative branch construction with a fixed mixing coefficient $\beta$, and use $r_{i}$ only for loss reweighting, which yields a stable supervised objective.
Reward Normalization.
For each pocket $p$, we sample $K$ molecules $\{M_{i}\}_{i=1}^{K}$ during rollout. We perform group-centering within each pocket:
$R_{\text{norm}}(M_{i},p):=R_{\text{raw}}(M_{i},p)-\frac{1}{K}\sum_{k=1}^{K}R_{\text{raw}}(M_{k},p).$
We then map the centered reward to an optimality probability $r_{i}\in[0,1]$ via
$r_{i}=0.5+0.5\cdot\mathrm{clip}\!\left(\frac{R_{\text{norm}}(M_{i},p)}{\max(Z,\epsilon)},-1,1\right),$
where $Z>0$ is a normalization scale (e.g., a running estimate of the global reward standard deviation) and $\epsilon$ is a small constant for numerical stability. Group-centering removes pocket-dependent reward bias while preserving within-pocket ranking signals.
Implicit Policy Parameterization.
We optimize on the same forward corruption processes as in pretraining. For each sampled molecule $M_{i}=(x_{1},v_{1})$, we sample $t\sim U[0,1]$ and corrupt both modalities:
$x_{t}\sim q_{t\mid 1}(x_{t}\mid x_{1}),\qquad v_{t}\sim\pi_{t}(v_{t}\mid v_{1}).$
The multi-flow model predicts clean coordinates via the $x$-prediction parameterization and atom-type logits:
$\hat{x}_{1,\theta}(x_{t},v_{t},p,t)$ $=x_{t}+(1-t)\,v_{\theta}(x_{t},v_{t},p,t),$
$\hat{\ell}_{\theta}(x_{t},v_{t},p,t)$ $\in\mathbb{R}^{N_{M}\times K}.$
We compute the corresponding anchor predictions $\hat{x}_{1,\theta_{\text{old}}}$ and $\hat{\ell}_{\theta_{\text{old}}}$ using the frozen anchor network.
Instead of learning separate guidance models, we implicitly construct positive/negative branches relative to the anchor policy. For continuous coordinates, define
$\hat{x}^{+}_{1,\theta}(x_{t},v_{t},p,t)=(1-\beta)\,\hat{x}_{1,\theta_{\text{old}}}(x_{t},v_{t},p,t)+\beta\,\hat{x}_{1,\theta}(x_{t},v_{t},p,t),$
$\hat{x}^{-}_{1,\theta}(x_{t},v_{t},p,t)=(1+\beta)\,\hat{x}_{1,\theta_{\text{old}}}(x_{t},v_{t},p,t)-\beta\,\hat{x}_{1,\theta}(x_{t},v_{t},p,t).$
For discrete atom types, linear interpolation in probability space is invalid, so we apply the same construction in logit (pre-softmax) space:
$\hat{\ell}^{\pm}_{\theta}(x_{t},v_{t},p,t)=(1\mp\beta)\,\hat{\ell}_{\theta_{\text{old}}}(x_{t},v_{t},p,t)\pm\beta\,\hat{\ell}_{\theta}(x_{t},v_{t},p,t).$
Intuitively, matching the positive branch to the ground truth pulls the model toward high-reward directions, while matching the negative branch encourages moving away from low-reward regions. This realizes negative-aware policy improvement without explicitly estimating $\pi^{+}$, $\pi^{-}$, or $\Delta$.
Joint Optimization Objective.
We optimize a reward-weighted supervised objective over the implicit positive and negative branches. Given $(M_{i},p)$, $t$, and $r_{i}$, we combine the continuous position loss $L_{\text{pos}}$ and the discrete cross-entropy loss $L_{\text{CE}}$:
(5) $L_{\text{NFT}}(\theta)=\mathbb{E}_{t,\,M_{i}}\Big[r_{i}\,\ell^{+}_{\text{pos}}+(1-r_{i})\,\ell^{-}_{\text{pos}}+r_{i}\,\ell^{+}_{\text{ce}}+(1-r_{i})\,\ell^{-}_{\text{ce}}\Big],$
where
$\ell^{\pm}_{\text{pos}}=L_{\text{pos}}\!\big(\hat{x}^{\pm}_{1,\theta}(x_{t},v_{t},p,t),\,x_{1}\big),\qquad\ell^{\pm}_{\text{ce}}=L_{\text{CE}}\!\big(\hat{\ell}^{\pm}_{\theta}(x_{t},v_{t},p,t),\,v_{1}\big).$
This objective remains fully supervised on the forward process while enabling policy improvement via reward-based reweighting, avoiding explicit likelihood-ratio estimation or policy-gradient updates.
3.6. Confidence Head
To explicitly estimate the quality of the generated molecules, we design an auxiliary Confidence Head module like PocketXMol (Peng et al., 2024). This optional module consists of two lightweight Multi-Layer Perceptrons (MLPs) that take the final invariant node embeddings from the backbone network as input. Specifically, one MLP predicts the confidence of atom types, formulated as a binary classification task to determine whether the predicted atom type matches the ground truth. The other MLP estimates structural reliability by regressing the spatial deviation between the generated and ground-truth coordinates. During the training phase, the confidence loss is integrated into the total objective as an auxiliary term, allowing the confidence head to be optimized jointly with the generative flow matching task.
4. Experiments
Model
Vina Score ($\downarrow$)
Vina Min ($\downarrow$)
Vina Dock ($\downarrow$)
Diversity ($\uparrow$)
QED ($\uparrow$) SA ($\uparrow$) Static Geometry ($\downarrow$) Clash ($\downarrow$)
Avg. Avg. Avg. Avg. Avg. Med. Avg. Med. JSDBL JSDBA Ratiocca Ratiocm
LiGAN -6.47 -7.14 -7.70 0.66 0.46 0.46 0.66 0.66 0.4645 0.5673 0.0096 0.0718
3DSBDD - -3.75 -6.45 0.70 0.48 0.48 0.63 0.63 0.5024 0.3904 0.2482 0.8683
GraphBP - - -4.57 0.79 0.44 0.44 0.64 0.64 0.5182 0.5645 0.8634 0.9974
Pocket2Mol -5.23 -6.03 -7.05 0.69 0.39 0.39 0.65 0.65 0.5433 0.4922 0.0576 0.4499
TargetDiff -5.71 -6.43 -7.41 0.72 0.49 0.49 0.60 0.60 0.2659 0.3769 0.0483 0.4920
DiffSBDD - -2.15 -5.53 - 0.49 0.49 0.34 0.34 0.3501 0.4588 0.1083 0.6578
DiffBP - - -7.34 - 0.47 0.47 0.59 0.59 0.3453 0.4621 0.0449 0.4077
FLAG - - -3.65 - 0.41 0.41 0.58 0.58 0.4215 0.4304 0.6777 0.9769
D3FG - -2.59 -6.78 - 0.49 0.49 0.66 0.66 0.3727 0.4700 0.2115 0.8571
DecompDiff -5.18 -6.04 -7.10 0.68 0.49 0.49 0.66 0.66 0.2576 0.3473 0.0462 0.5248
MolCraft -6.15 -6.99 -7.79 0.72 0.48 0.48 0.66 0.66 0.2250 0.2683 0.0264 0.2691
VoxBind -6.16 -6.82 -7.68 - 0.54 0.54 0.65 0.65 0.2701 0.3771 0.0103 0.1890
MolFORM -5.42 -6.42 -7.50 0.78 0.48 0.49 0.60 0.58 0.3225 0.5535 0.0310 0.4474
TAGMol -7.02 -7.95 -8.59 0.63 0.55 0.56 0.56 0.55 0.2389 0.5015 0.0237 0.3190
DecompOpt -5.75 -6.58 -7.63 0.69 0.48 0.45 0.65 0.65 - - - -
MolJO -7.52 -8.33 -9.05 0.66 0.56 0.57 0.78 0.77 0.4287 0.4555 0.0240 0.2696
Alidiff -7.07 -8.09 -8.90 0.73 0.50 0.50 0.57 0.56 0.3418 0.5333 0.0268 0.3324
MolFORM-DPO -6.16 -7.18 -8.13 0.77 0.50 0.51 0.65 0.63 0.3215 0.5584 0.0188 0.2525
MolFORM-RL -7.60 -8.37 -9.24 0.75 0.50 0.51 0.68 0.67 0.6098 0.4430 0.0331 0.3814
Reference -6.36 -6.71 -7.45 - 0.48 0.47 0.73 0.74 - - - -
Table 1. Combined results for binding affinity, chemical properties, and geometry/clash metrics. (↑)/(↓) denote better. Top 2 results are marked in bold and underlined. The result from baseline model is quoted from  Lin et al. (2024).
4.1. Experiment Setup
Datasets
Our experiments were conducted using the CrossDocked2020 dataset (Francoeur et al., 2020). Consistent with prior studies  (e.g., Peng et al., 2022; Guan et al., 2023, 2024), we adhered to the same dataset filtering and partitioning methodologies. We further refined the 22.5 million docked protein binding complexes, characterized by an RMSD $1\AA$ , and sequence identity less than 30%. This resulted in a dataset comprising 100,000 protein-binding complexes for training, alongside a set of 100 novel complexes designated for testing.
DPO dataset
Our data processing strategy for DPO follows the methodology proposed in  Gu et al. (2024). We preprocess the dataset into a preference format $\mathcal{D}=\{(\mathbf{p},\mathbf{m}^{w},\mathbf{m}^{l})\}$, where $\mathbf{p}$ denotes the protein pocket, $\mathbf{m}^{w}$ the preferred ligand, and $\mathbf{m}^{l}$ the less preferred one. For each pocket, we sample two candidate ligands and assign preference based on a user-defined reward, mainly binding energy (e.g., Vina score). Since the affinity labels are continuous, we follow the strategy in Gu et al. (2024) and choose the molecule with the worst score as the dispreferred sample $\mathbf{m}^{l}$, which encourages a larger reward gap between $\mathbf{m}^{w}$ and $\mathbf{m}^{l}$.
RL dataset
Distinct from static offline datasets, we adopt an on-policy data collection strategy. Using the processed CrossDocked2020 pockets as the environment, we dynamically generate candidate ligands via policy rollouts during each training iteration. Valid molecules are evaluated using specified reward functions (e.g., Vina, QED, SA). To stabilize training, rewards are normalized within each pocket—centered, scaled, clipped to [1,1][-1,1], and mapped to [0,1][0,1]—forming temporary tuples used exclusively for the current parameter update.
RL rollout and scoring.
We maintain an EMA anchor policy (decay $0.995$) for sampling and as the reference network. For each update, we sample B=4B{=}4 pockets and generate K=8K{=}8 ligand candidates per pocket using a 100-step log-time multi-flow sampler (atom-count prior; discrete temperature $0.01$; discrete noise $1.0$). We evaluate each generated ligand by AutoDock Vina and compute a normalized synthetic accessibility (SA) score. Given tuples $(p,m_{p,i},r_{p,i})$, we sample $t\sim\mathcal{U}(0,1)$ and apply the same forward corruptions as pretraining. We form implicit positive/negative branches by mixing current and anchor predictions with $\beta=\beta_{\mathrm{discrete}}=0.3$, and minimize the reward-weighted DiffusionNFT objective using Adam (lr $5\times 10^{-7}$) with gradient clipping $8.0$.
Model architecture
Inspired from recent progress in equivariant neural networks (Satorras et al., 2021), we model the interaction between the ligand molecule atoms and the protein atoms with a SE(3)-Equivariant GNN, the atom hidden embedding and coordinates are updated alternately in each layer, which follows Guan et al. (2023). Our model architecture is plotted in Figure 1.
Baselines.
We select several baseline models from CBGbench (Lin et al., 2024) for comparison with our method. Early structure-based drug design (SBDD) methods are built on voxel grids with deep neural networks, such as LiGAN (Ragoza et al., 2022), which generates atom voxelized density maps using variational autoencoders (VAE) and convolutional neural networks (CNNs), and 3DSBDD (Luo et al., 2022), which predicts atom types on grids with graph neural networks (GNNs) in an auto-regressive manner. The development of equivariant graph neural networks (EGNNs) enables direct generation of 3D atom positions, as seen in Pocket2Mol (Peng et al., 2022) and GraphBP (Liu et al., 2022a), which use auto-regressive strategies with normalizing flows. Diffusion-based methods such as TargetDiff (Guan et al., 2023), DiffBP (Lin et al., 2022), and DiffSBDD (Schneuing et al., 2024) generate atom types and positions using denoising diffusion probabilistic models. Recent methods incorporate domain knowledge to guide generation: FLAG (Zhang et al., 2023) and D3FG (Lin et al., 2023) use fragment motifs for coarse molecular generation; DecompDiff (Guan et al., 2024) uses scaffold and arm clustering with Gaussian process models for atom positions. More recent advances like MolCraft (Qu et al., 2024) and VoxBind (Pinheiro et al., 2024) apply new generative modeling strategies, including Bayesian flow networks and voxel-based diffusion with walk-jump sampling.
We additionally include three recent strong baselines: TAGMol (Dorna et al., 2024), which uses target-aware gradient guidance to steer diffusion sampling toward better affinity and auxiliary properties; DecompOpt (Zhou et al., 2024a), a controllable decomposed diffusion framework for structure-based molecular optimization; and MolJO (Qiu et al., 2024), a gradient-guided Bayesian-update approach that jointly optimizes discrete atom types and continuous coordinates.
Figure 3. Comparing the distribution for distances of allatom (top row) and carbon-carbon pairs (bottom row) for reference molecules in the test set (gray) and model generated molecules (color).
Figure 4. Visualizations of reference molecules and generated ligands for protein pockets (4yhj) generated by Reference, Targetdiff, MolFORM, MolFORM-DPO and MolFORM-RL. Vina score, QED, and SA are reported below.
4.2. Experiment result
Evaluation
We collect all generated molecules across 100 test proteins and evaluate generated ligands from three aspects: binding affinity, molecular properties and molecular structures. For target binding affinity and molecular properties, we present our results under the best setting as MolFORM. Following previous work, we utilize AutoDock Vina (Eberhardt et al., 2021) for binding affinity estimation. For binding affinity, we report the Vina Score, which evaluates the initially generated binding pose; the Vina Min, obtained after local energy minimization; and the Vina Dock, representing the lowest energy score from a global re-docking procedure using grid-based search. For molecular properties, we primarily report QED (Quantitative Estimate of Drug-likeness) and SA (Synthetic Accessibility). These results are summarized in Table 1. For molecular structures, we first evaluate several geometry-related properties following the setup in  Lin et al. (2024). These include (i) JSDBL and (ii) JSDBA, which measure the divergence in bond length and bond angle distributions between generated and reference molecules, reflecting structural realism. (iii) Ratiocca denotes the proportion of atoms with steric clashes—defined as van der Waals overlaps $\geq 0.4$Å—with protein atoms. (iv) Ratiocm captures the fraction of generated molecules that contain any such clashes. These results are also included in Table 1. Further structural evaluations, including Root Mean Square Deviation (RMSD), ring size distributions, and substructure-level bond length JSD, are detailed in the Appendix (Figure 7, Table 3, and Table 4).
Figure 5. Effectiveness of the Confidence Head. Selecting a smaller ratio of top-confidence samples yields significantly better (lower) Vina scores.
Figure 6. Median Vina energy for different generated molecules (TargetDiff vs. MolCraft vs. MolJO vs. MolFORM RL) across 100 testing binding targets. Binding targets are sorted by the median Vina energy of generated molecules. Lower Vina energy means a higher estimated binding affinity.
Result analysis
As summarized in Table 1, our base model MolFORM demonstrates strong generative capabilities. While achieving comparable binding affinity to the diffusion baseline TargetDiff, MolFORM significantly outperforms it in structural validity and diversity. Specifically, MolFORM reduces the steric clash ratio (Ratiocca) from 0.0483 to 0.0310 and achieves a higher diversity score (0.78 vs. 0.72), indicating its ability to explore a broader chemical space with physically plausible structures.The introduction of preference alignment yields substantial improvements. MolFORM-DPO excels in balancing generation quality and diversity. It achieves a remarkable reduction in steric clashes (0.0188) and improves the Vina Score to -6.16, while maintaining a high diversity of 0.77.
Furthermore, MolFORM-RL achieves state-of-the-art performance in binding affinity, with a Vina Score of -7.60, surpassing recent strong baselines such as TAGMol (-7.02) and MolJO (-7.52). This suggests that our alignment strategies successfully steer the generative flow toward chemically favorable manifolds, preserving the intrinsic chemical priors learned during pre-training. Meanwhile, MolFORM-RL maintains a high diversity score of 0.75, highlighting a key distinction from other fine-tuning approaches that often improve affinity at the cost of diversity.
Finally, we validate the effectiveness of the auxiliary Confidence Head. As illustrated in Figure 5, the predicted confidence scores exhibit a strong correlation with the ground-truth structural quality (measured by RMSD). This indicates that the confidence head serves as a reliable estimator, enabling efficient ranking and filtration of high-fidelity candidates.
4.3. Ablation study on Reward design
We employed three distinct reward formulations. The first one is formulated as a linear combination of normalized Vina Score and SA as follows:
[…]
where $\text{Clip}(\cdot,\cdot,\cdot)$ denotes the clipping operation. The second type combines QED and Vina score, while the third corresponds to QED and SA. We observed that the first two types yield sustained performance gains; however, the third type does not lead to significant improvements in the Vina score. Detailed reward curves illustrating the online RL training process are provided in the Figure 8.
Table 2. Ablation study on different reward formulations. Reward 1 (Vina + SA) achieves the best binding affinity, while Reward 2 improves QED. Reward 3 shows limited improvement on Vina score.
Metric Reward 1 Reward 2 Reward 3
(Vina + SA) (QED + Vina) (QED + SA)
Vina Score ($\downarrow$) -7.60 -7.08 -4.38
Vina Min ($\downarrow$) -8.37 -8.13 -5.84
QED ($\uparrow$) 0.50 0.60 0.56
SA ($\uparrow$) 0.68 0.56 0.66
5. Conclusions
In this work, we introduce MolFORM, a multimodal flow matching framework for protein-specific molecular generation that jointly models discrete atom types and continuous 3D coordinates. We further show that online reinforcement learning provides a powerful mechanism for aligning flow-based generative models with biochemical objectives. On the CrossDocked2020 benchmark, MolFORM-RL achieves state-of-the-art binding affinity. Moreover, our reinforcement learning framework holds strong potential for extension to other structure-based drug design (SBDD) generative models.
Appendix A Algorithm
Algorithm 1 Forward-Process Negative-aware Online Fine-tuning for Conditional Diffusion-based Ligand Generation
1:Pretrained conditional diffusion/flow policy $v_{\mathrm{ref}}$; pocket dataset $\mathcal{C}$ (conditions $c$); raw reward function $r_{\mathrm{raw}}(\mathbf{x}_{0},c)\in\mathbb{R}$; rollout size $K$; mixing coefficient $\beta_{\mathrm{NFT}}$; learning rate $\lambda$; EMA update ratio $\eta_{i}$; diffusion schedule / forward corruption operator $q_{t|0}$.
2:Notation: a ligand sample is $\mathbf{x}_{0}=(\mathbf{s}_{0},\mathbf{r}_{0})$ (atom types $\mathbf{s}_{0}$ and 3D coordinates $\mathbf{r}_{0}$).
3:Initialize: sampling policy $v_{\mathrm{old}}\leftarrow v_{\mathrm{ref}}$, training policy $v_{\theta}\leftarrow v_{\mathrm{ref}}$, replay buffer $\mathcal{D}\leftarrow\emptyset$.
4:for iteration $i=1,2,\dots$ do
5:  for each sampled pocket condition $c\sim\mathcal{C}$ do $\triangleright$ Rollout / Data Collection
6:   Sample $K$ ligand structures $\{\mathbf{x}_{0}^{k}\}_{k=1}^{K}\sim\pi_{\mathrm{old}}(\cdot\,|\,c)$ using any black-box solver.
7:   Compute rewards $r_{\mathrm{raw}}^{k}\leftarrow r_{\mathrm{raw}}(\mathbf{x}_{0}^{k},c)$ for $k=1..K$.
8:   Group-normalize rewards: $r_{\mathrm{norm}}^{k}\leftarrow r_{\mathrm{raw}}^{k}-\frac{1}{K}\sum_{j=1}^{K}r_{\mathrm{raw}}^{j}$.
9:   Convert to optimality probability $r^{k}\in[0,1]$:
$r^{k}\leftarrow\frac{1}{2}+\frac{1}{2}\,\mathrm{clip}\!\left(\frac{r_{\mathrm{norm}}^{k}}{Z_{c}},-1,1\right),$
where $Z_{c}>0$ is a normalizer (e.g., global reward std).
10:   Add tuples to buffer: $\mathcal{D}\leftarrow\mathcal{D}\cup\{(c,\mathbf{x}_{0}^{k},r^{k})\}_{k=1}^{K}$.
11:  end for
12:  for each minibatch $\{(c,\mathbf{x}_{0},r)\}\subset\mathcal{D}$ do $\triangleright$ Policy Optimization on Forward Process
13:   Sample time $t\sim\mathcal{U}(0,1)$ and noises (Gaussian for coordinates; discrete corruption noise for atom types).
14:   Forward corruption: sample noisy state $\mathbf{x}_{t}\sim q_{t|0}(\cdot\,|\,\mathbf{x}_{0})$.
15:   Compute the standard diffusion/flow regression target $\mathbf{v}$ under your parameterization (e.g., velocity/flow/score target).
16:   Implicit positive branch:
$v_{\theta}^{+}(\mathbf{x}_{t},c,t)\leftarrow(1-\beta_{\mathrm{NFT}})\,v_{\mathrm{old}}(\mathbf{x}_{t},c,t)+\beta_{\mathrm{NFT}}\,v_{\theta}(\mathbf{x}_{t},c,t).$
17:   Implicit negative branch:
$v_{\theta}^{-}(\mathbf{x}_{t},c,t)\leftarrow(1+\beta_{\mathrm{NFT}})\,v_{\mathrm{old}}(\mathbf{x}_{t},c,t)-\beta_{\mathrm{NFT}}\,v_{\theta}(\mathbf{x}_{t},c,t).$
18:   Update parameters by minimizing the negative-aware forward loss:
$\theta\leftarrow\theta-\lambda\nabla_{\theta}\Big(r\|v_{\theta}^{+}-\mathbf{v}\|_{2}^{2}+(1-r)\|v_{\theta}^{-}-\mathbf{v}\|_{2}^{2}\Big).$
19:  end for
20:  $\triangleright$ Online EMA update of sampling policy (off-policy)
21:  $\theta_{\mathrm{old}}\leftarrow\eta_{i}\,\theta_{\mathrm{old}}+(1-\eta_{i})\,\theta$;  $\mathcal{D}\leftarrow\emptyset$.
22:end for
23:Output: fine-tuned conditional policy $v_{\theta}$.
Appendix B Additional Experimental Results
In this section, we provide detailed structural and chemical assessments of the generated molecules to validate their geometric integrity and distribution consistency.
Structural Deviation Analysis
Figure 7 illustrates the Median Root Mean Square Deviation (RMSD) for rigid fragments before and after force-field optimization. RMSD serves as a standard metric to evaluate structural deviation; lower values indicate higher structural consistency. The results reflect the extent to which rigid fragments preserve their geometric integrity, demonstrating that our generated structures are chemically stable and require minimal adjustment during optimization.
Chemical Property Distributions
We further evaluate the chemical validity of the generated molecules through bond length distributions and ring size analysis. Table 3 presents the Jensen-Shannon Divergence (JSD) of bond lengths for various bond types compared to the reference dataset. Lower JSD scores indicate that the generated bond lengths closely match the ground truth distribution. Additionally, ring size distributions are detailed in Table 4, confirming that the model generates chemically plausible ring structures.
Figure 7. Median RMSD for rigid fragment before and after the force-field optimization.
Table 3. JSD Bond Length Comparisons across different methods.
Method C-C C-N C-O C=C C=N C=O
LiGan 0.4986 0.4146 0.4560 0.4807 0.4776 0.4595
3DSBDD 0.2090 0.4258 0.5478 0.5170 0.6701 0.6448
GraphBP 0.5038 0.4231 0.4973 0.6235 0.4629 0.5986
Pocket2Mol 0.5667 0.5698 0.5433 0.4787 0.5989 0.5025
TargetDiff 0.3101 0.2490 0.3072 0.1715 0.1944 0.3629
DiffSBDD 0.3841 0.3708 0.3291 0.3043 0.3473 0.3647
DiffBP 0.5704 0.5256 0.5090 0.6161 0.6314 0.5296
FLAG 0.3460 0.3770 0.4433 0.4872 0.4464 0.4292
D3GF 0.4244 0.3227 0.3895 0.3860 0.3570 0.3566
DecompDiff 0.2562 0.2007 0.2361 0.2590 0.2844 0.3091
MolCraft 0.2473 0.1732 0.2341 0.3040 0.1459 0.2250
VoxBind 0.3335 0.2577 0.3507 0.1991 0.1459 0.3334
MolFORM 0.4852 0.3876 0.3830 0.3919 0.5496 0.4372
Table 4. Proportion (%) of different ring sizes in reference and generated ring-structured molecules, where 3-Ring denotes three-membered rings and the like.
3-Ring 4-Ring 5-Ring 6-Ring
Reference 4.0 0.0 49.0 84.0
Train 3.8 0.6 56.1 90.9
AR 50.8 0.8 35.8 71.9
Pocket2Mol 0.3 0.1 38.0 88.6
FLAG 3.1 0.0 39.9 84.7
TargetDiff 0.0 7.3 57.0 76.1
DecompDiff 9.0 11.4 64.0 83.3
IPDiff 0.0 6.4 51.0 83.7
MolCRAFT 0.0 0.6 47.0 85.1
DecompOpt 6.8 11.8 61.4 89.8
TAGMol 0.0 8.5 62.5 82.6
MolJO 0.0 0.3 44.4 97.6
MolFORM 0.0 3.2 52.0 84.4
Ring size
Furthermore, we analyze the ring size distribution to assess the topological realism of the generated molecules. Table 4 compares the frequencies of different ring sizes (e.g., 3-, 4-, 5-, and 6-membered rings) against the reference dataset. The results demonstrate that our model generates chemically plausible ring structures, avoiding the over-generation of unstable small rings or unrealistic large fused systems often observed in baseline methods.
Chamfer loss Impact
We conducted ablation studies on the Chamfer DPO loss component in Table 5. Our experiments demonstrate that incorporating the Chamfer distance significantly enhances the model’s ability to preserve molecular geometric fidelity.
Table 5. Ablation study: Comparison of model performance with (Chamfer-w) and without (Chamfer-wo) Chamfer loss in Vanilla Multi-model flow matching.
Metric Chamfer-w Chamfer-wo
QED ($\uparrow$) 0.48 0.41
SA ($\uparrow$) 0.60 0.59
Vina Score ($\downarrow$) -5.42 -4.67
Vina Min ($\downarrow$) -6.42 -5.76
JSDBL ($\downarrow$) 0.3225 0.4546
JSDBA ($\downarrow$) 0.5535 0.6173
Abalation on DPO training
We re-trained a version of Targetdiff and applied DPO-based fine-tuning on the model. We observed that applying DPO to Targetdiff led to improvements of 2%, 2%, 2% and 3% in QED, SA, Vina score, and Vina min, while MolFORM achieved improvements of 4%, 8%, 14% and 12% on the same metrics in Table 6. This suggests that our model has greater potential to benefit from DPO fine-tuning. MultiFlow explicitly separates the generation of discrete (e.g., atom types) and continuous (e.g., 3D coordinates) modalities via dedicated flow-based branches. This factorized structure enables DPO to assign fine-grained preference signals to each modality during optimization. As DPO compares generation quality between molecule pairs, the modular design of MultiFlow allows more targeted updates—for example, refining atom types without perturbing geometry, or vice versa. This structural disentanglement enhances the model’s responsiveness to reward signals, reduces gradient interference, and ultimately makes preference optimization more effective and stable.
Table 6. Comparison of model performance before and after DPO fine-tuning. The reported numbers are average values here. The results of vanilla TargetDiff are obtained by retraining the model.
Metric TargetDiff TargetDiff-DPO MolFORM MolFORM-DPO
Vina Score ($\downarrow$) -5.47 -5.58 -5.42 -6.16
Vina Min ($\downarrow$) -6.39 -6.59 -6.42 -7.18
QED ($\uparrow$) 0.46 0.47 0.48 0.50
SA ($\uparrow$) 0.60 0.61 0.60 0.65
Training detail
To ensure training stability during the alignment phase, we employed significantly lower learning rates compared to the pre-training stage. Specifically, the learning rate was set to $5\times 10^{-8}$ for DPO and $5\times 10^{-6}$ for online RL, which are considerably smaller than the base model’s learning rate of $5\times 10^{-5}$. For the MolFORM-RL configuration, we set the KL regularization coefficients $\beta$ and $\beta_{\text{discrete}}$ to 0.3 to prevent excessive deviation from the prior distribution. Additionally, we utilized an Exponential Moving Average (EMA) with a decay rate of 0.995 for target network updates and applied gradient clipping with a maximum norm of 8.0 to further stabilize the optimization process. Detailed reward curves illustrating the training stability are provided in the Appendix. We observed that employing a composite reward of QED, SA, and Vina score yields superior performance, effectively enhancing binding affinity while maintaining high drug-likeness.
Figure 8. Online RL reward curves for Reward 1 and Reward 2.